# Matching Markets Meet LLMs: Algorithmic Reasoning with Ranked Preferences

### v1

@@ 3 Generating Stable Solutions
We consider two sub-categories for generating matchings depending on declarative knowledge about algorithms: i) prompt without specifying any algorithm, and ii) prompt with exact step-by-step instructions of how to execute the DA algorithm [27] (see Section 2.1 for details). Our ablation studies showed that the above prompting strategies did not result in qualitatively different outcomes, as all models were able to correctly identify the requirement for considering preferences, the DA algorithm, and its execution steps. [^3] Furthermore, converting the stable-marriage setting to a task-scheduling setting [24], where “men” and “women” are replaced by “workers” and “tasks” (respectively), does not have a significant impact on performance. The detailed results are presented in Appendix D.
@@
Difficulty, Model Size, and Reasoning. Figure 1 demonstrates the performance of the models in generating stable outcomes. Baseline models without explicit reasoning mechanisms are unable to solve even Easy instances, whereas reasoning-enabled models achieve high accuracy on Easy instances but suffer dramatic performance drops on Hard instances. Furthermore, for Hard problems, even reasoning models frequently produce invalid outputs or fail to return any solution. Interestingly, Qwen-QwQ-32B significantly outperforms a much larger model, DeepSeek-70B, indicating that LLMs’ combinatorial reasoning capability does not necessarily scale with model size. [^4] Throughout the paper, all statistical comparisons between the percentages of stable solutions returned (across LLMs or across treatments) are done using Fisher’s Exact test [25]. Similarly, any two distributions of Instability or Optimality Rate are statistically compared using Welch’s T-test [71].
IC vs. ML Profiles. Under Impartial Culture (IC) profiles, the number of stable solutions can grow exponentially as the problem scales (increase of $n$) [44]. This combinatorial explosion poses a significant challenge for LLMs attempting to identify stable solutions, especially when solely using implicit reasoning over preference lists (without executing a concrete matching algorithm). In contrast, master-list profiles (ML)—irrespective of the underlying sampling method used to generate preferences—admit exactly one stable solution. Moreover, this unique stable matching can be constructed in $\mathcal{O}(n)$ steps by (i) extracting the common master list and then (ii) pairing agents in order of their shared priority [37]. See the details of the algorithm in Appendix C.
We observe that there is a significant performance gap between ML and IC instances—this disparity is especially marked for DeepSeek-R1. With IC profiles and Hard instances, all models are unable to compute a stable solution. [^5] Gemini-2.5-Pro is the only model with a positive success rate ($=8\%$) with IC preferences Hard instances. They perform slightly better under ML profiles, and while this performance drops for Hard instances, these models almost never return invalid or partial matchings.
Prompting Techniques. Prompt-engineering techniques have been empirically demonstrated to enhance the performance of LLMs on mathematical reasoning and formal logic inference tasks [70, 75, 3]. We evaluated a range of prompt-engineering techniques—including few-shot prompting and Chain-of-Thought (CoT) prompting, which supply exemplar “thought processes” and intermediate reasoning steps—in an attempt to bolster LLM performance. However, none of these strategies qualitatively improved on medium- or hard-difficulty instances. See Appendix D for further details.
@@ 3.1 Measuring Instability
A natural question is how far LLM-produced responses deviate from stable outcomes. To quantify this, we use two complementary metrics: instability rate and optimality rates (see Section 2.2). The instability rate directly reflects the distance from any stable solution, whereas the optimality rate implicitly evaluates the model’s success in executing the underlying matching procedure. Figure 3 and Figure 3 illustrate comparisons of LLMs with a baseline of randomly generated outcomes. [^6] Note that the plots only illustrate unstable but valid one-to-one outputs.
@@
Interestingly, the performance distinction between basic and reasoning models becomes less clear. While the intermediate reasoning models return a lower instability rate in Easy and Medium problems, their performance significantly drops in larger-scale problems (Hard). In fact, the performance of DeepSeek-70B becomes worse than even basic non-reasoning models. We attribute this behavior to the model’s diminished capacity for handling larger input lengths, a hypothesis supported by their lower proportion of valid outcomes (as seen in Figure 1).
@@ 4 Resolving Instability
Generating stable solutions requires both exact reasoning over agents’ preference lists and the execution of a stability-guaranteeing procedure (e.g. the DA algorithm). As demonstrated in Section 3, all evaluated models—irrespective of their reasoning capabilities—exhibit severe performance degradation as the problem size grows. This leads to the natural question of whether these models can resolve instability in a given matching–a task that entails detecting blocking pairs through preference reasoning and applying an appropriate sequence of adjustments to restore stability.
We provide LLMs with unstable (but valid) matchings along with preference profiles, and instruct them to convert these initialized solutions to stable matchings. To assess how the instability rate may influence solution quality, we distinguish two classes of initial matchings: (i) One-BP, matchings containing exactly one blocking pair (i.e., “almost stable”) such that their stability may be resolved through simpler proposal-rejection iterations, and (ii) Random, matchings sampled uniformly at random from the set of all valid one-to-one pairings, which typically contain a large number of blocking pairs and thus exhibit high degrees of instability. See Appendix E for detailed steps and pseudo-code for generating one-BP and random initialization. Note that starting from an arbitrary matching, sequentially resolving blocking pairs may result in a cycle—as shown by Knuth [43]. However, a random sequence converges to stability with probability one [60, 1].
@@
Table 1 displays the fraction of responses in which LLMs return stable matchings when asked to resolve the above types of unstable matching. Surprisingly, our experiments illustrate that in the task of resolving instability, the performance of all evaluated models does not exceed—and even degrades—their performance in generating stable solutions. This behavior persists regardless of initial matchings (One-BP or Random) and LLMs’ reasoning capability. In fact, on Hard instances, the output returned by advanced reasoning models on One-BP matchings (i.e., containing a single blocking pair) contains substantially more than one blocking pair. In other words, even for the most basic instances, LLMs often introduce additional instabilities beyond the original single violation. We elaborate on this in Appendix E, demonstrating how all models, including those with advanced reasoning, often return solutions with a higher instability rate, highlighting their inability to systematically eliminate blocking pairs in accordance with preference lists.
@@ 5 Detecting Instability
The findings in previous sections raise the question of whether LLMs can reliably detect instability in a given matching—a simpler task that involves iterating over each unmatched pair to determine whether both agents prefer one another over their assigned partners. This procedure requires only a straightforward comparison of preferences and requires $\mathcal{O}(n^{2})$ steps.
@@
Reasoning Models: Preferences and Blocking Pairs. Figure 4 reveals an interesting observation about reasoning models: their performance is influenced by the number of blocking pairs in the matching being evaluated—similar to the observations in Section 4. They achieve a high accuracy on identifying random matchings (which have a larger number of blocking pairs) as unstable and a significantly lower accuracy with matchings that have at most one blocking pair.
Basic Models and Hallucination. Interestingly, the non-reasoning models achieve a high accuracy ($80\%$) with both types of unstable matchings, and extremely low ($20\%$) accuracy with stable matchings. Note that the performance across all models is qualitatively similar in ML and IC profiles, even though each ML profile admits a unique stable solution (thus, identical reports for Optimal and Pessimal). See Appendix F for further analysis. A manual analysis of non-reasoning models uncovers frequent hallucinations about blocking pairs, resulting in a systematic bias toward classifying matchings as unstable. This can be largely attributed to misinterpretations of the input preferences.
@@ 6 Reasoning about Ranked Preferences
Many advanced reasoning paradigms, ranging from causal inference [12] and counterfactual analysis to game-theoretic decision making, depend fundamentally on the ability to compare and evaluate alternative choices. As demonstrated thus far, even advanced reasoning models often fail to execute the step-by-step procedures of combinatorial algorithms when those procedures operate over ranked preference lists. This shortcoming motivates the question of whether current LLMs can truly reason about preferences, as opposed to merely applying preferences in generating responses heuristically.
@@
Hierarchical, level-wise reasoning evaluations have been proposed recently in domains such as causal inference of LLMs [12]. For example, an L1 question is “Who is agent W5’s, 4th-most preferred agent?”, and an L2 question ‘Would agent W5, prefer M8 over M7?”
@@
Basic models have low accuracy in all levels of difficulty even on small instances, which is probably the reason behind their inability to compute or detect stable solutions (as discussed in Section 3 and Section 5). In basic models (e.g. Llama-3.3-70B) and (non-advanced) reasoning models (e.g. DeepSeek-70B), the size of the problem has a greater impact on accuracy as compared to the level of the question, indicating their difficulty of handling larger inputs. Although advanced-reasoning models have significantly higher accuracy rates compared to other models, they still make minor errors, especially in larger profiles (Hard). Generating stable solutions for these instances often requires a larger number of reasoning steps over many preference lists, causing minor errors to compound (small errors multiply!).
@@ 7 Performance Improvement through LoRA Fine-Tuning
Supervised fine-tuning has proven to be effective in enhancing the reasoning capabilities of LLMs in a variety of tasks, including mathematical problem solving [13, 46], logical reasoning [57], and code generation [51]. Markeeva et al. [54] demonstrate how fine-tuning a small LLM (2B parameters) can significantly improve LLMs’ ability to execute textbook algorithms (e.g., sorting an array, finding the shortest path between two nodes on a graph, etc.). Hence, we evaluate whether fine-tuning can be used to enhance LLMs’ ability to compute stable matchings—a task requiring reasoning over ranked preferences in addition to algorithmic understanding. To this end, we perform LoRA fine-tuning [34] on three reasoning models, including Qwen-QwQ-32B and two smaller models, DeepSeek-8B and DeepSeek-14B, for Generating Stable Solutions task. Additionally, we also evaluate whether fine-tuning can mitigate errors made by LLMs in the Preference Reasoning task (Section 6).
Training. Let $\mathcal{D}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N}$ be the fine-tuning dataset for a given task. Each pair consists of an input instance $\mathbf{x}^{(i)}$ and the desired model completion $\mathbf{y}^{(i)}$. The input instance $\mathbf{x}^{(i)}$ is made up of four components: (i) a generic system-prompt $s$, (ii) a high-level instruction $u$, (iii) the preference profile $p^{(i)}$, and (iv) the task-prompt $t^{(i)}$. The desired completion $\mathbf{y}^{(i)}$ consists of two components, (i) a chain-of-thought reasoning trace $r^{(i)}$, and (ii) the answer $a^{(i)}$ in the desired format. Each model is fine-tuned with standard next-token cross-entropy on the concatenated sequence $\mathbf{z}^{(i)}=\mathbf{x}^{(i)}||\mathbf{y}^{(i)}$. We separately fine-tune each model for the Generating Stable solutions task ($N=10,000$) and the Preference Reasoning task ($N=9,000$). See Appendix G for details of the fine-tuning setup.
Improvement. Fine-tuning LLMs with data containing synthetically generated reasoning traces substantially improves their performance on both tasks, as evidenced in Table 2.
@@
In fact, this approach enhances the performance of Qwen-QwQ-32B, a (non-advanced) reasoning model, to a success rate of $100\%$ in computing stable matchings with both Easy and Medium instances, significantly outperforming advanced-reasoning models. Fine-tuning also clearly improves smaller models, i.e. DeepSeek-8B and DeepSeek-14B, both in terms of achieving stability and Instability Rate. [^7] Interestingly, this improvement is clearer for ML instances, where both models achieve a $100\%$ success rate at the Easy level and $>80\%$ success rate at the Medium level. Similar results are obtained for the Preference Reasoning tasks, with the error-rate reducing to 00 at the Easy and Medium levels.  [^8] The only exception being DeepSeek-8B at the Medium level.
@@
[…]
@@ 8 Concluding Remarks
We summarize the performance of LLMs across all four tasks in Table 3, reflecting the clear hierarchy between advanced-reasoning models, (non-advanced) reasoning models, and basic models. The limitations in reliably reasoning about ranked preferences raise concerns about the viability of LLMs as agents acting on behalf of users in market-oriented or preference-sensitive decision-making settings, limits their capacity to negotiate complex user preferences, and hinders efforts in developing pluralistic techniques (e.g., constitutional AI [2] and social choice-theoretic [15]) for value alignment that are inherently based on aggregating rankings.
Open-Source vs. Closed-Source models. Among the models that we evaluate, Gemini-2.5-Pro (a closed-source model) emerges as the most capable across all tasks. While DeepSeek-R1 (open-source) broadly outperforms OpenAI’s o3-mini (closed-source), it performs much worse with IC instances than with ML instances. While both basic models struggle on all tasks, Gemini-2.0-Flash (closed-source) marginally outperforms Llama-3.3-70B (open-source) on various metrics. Given the promising improvement yielded by fine-tuning an open-source reasoning model, i.e. Qwen-QwQ-32B, it is worth exploring strategies also that enable it to handle large inputs.
Beyond Linear Preferences. Our current evaluation paradigm considers complete and strict linear preferences. In real-world scenarios, however, preferences involve complexities such as incompleteness, indifference between alternatives, and capacity constraints [18, 45, 53]. As a result, algorithms to compute stable solutions in such settings are far more complex and solutions are often intractable [52]. A meaningful next step would be to examine how AI models respond to such situations.

### v2

@@ 3 Generating Stable Solutions
We consider two sub-categories for generating matchings depending on declarative knowledge about algorithms: i) prompt without specifying any algorithm, and ii) prompt with exact step-by-step instructions on how to execute the DA algorithm [gale1962college] (see Section˜2.1 for details). Our ablation studies showed that the above prompting strategies did not result in qualitatively different outcomes, as all models were able to correctly identify the requirement for considering preferences, the DA algorithm, and its execution steps. [^2] Furthermore, converting the stable-marriage setting to a task-scheduling setting [Fish2025EconEvals], where “men” and “women” are replaced by “workers” and “tasks” (respectively), does not have a significant impact on performance. The detailed results are presented in Appendix˜D.
@@
Difficulty, Model Size, and Reasoning. Figure˜2 demonstrates the performance of the models in generating stable outcomes. Baseline models without explicit reasoning mechanisms are unable to solve even Easy instances, whereas reasoning-enabled models achieve high accuracy on Easy instances but suffer dramatic performance drops on Hard instances. Furthermore, for Hard problems, even reasoning models frequently produce invalid outputs or fail to return any solution. Interestingly, Qwen-QwQ-32B significantly outperforms a much larger model, DeepSeek-70B, indicating that LLMs’ combinatorial reasoning capability does not necessarily scale with model size. [^3] Throughout the paper, all statistical comparisons between the percentages of stable solutions returned (across LLMs or across treatments) are done using Fisher’s Exact test [FishersExact]. Similarly, any two distributions of Instability or Optimality Rate are statistically compared using Welch’s T-test [WelchTTest].
IC vs. ML Profiles. Under Impartial Culture (IC) profiles, the number of stable solutions can grow exponentially as the problem scales (increase of $n$) [knuth1997stable]. This combinatorial explosion poses a significant challenge for LLMs attempting to identify stable solutions, especially when solely using implicit reasoning over preference lists (without executing a concrete matching algorithm). In contrast, master‐list profiles (ML)—irrespective of the underlying sampling method used to generate preferences—admit exactly one stable solution. Moreover, this unique stable matching can be constructed in $\mathcal{O}(n)$ steps by (i) extracting the common Master-list and then (ii) pairing agents in order of their shared priority [irving2008stable]. See the details of the algorithm in Appendix˜C.
We observe that there is a significant performance gap between ML and IC instances—this disparity is especially marked for DeepSeek-R1. With IC profiles and Hard instances, all models are unable to compute a stable solution. [^4] Gemini-2.5-Pro is the only model with a positive success rate ($8\%$) with IC preferences Hard instances. They perform slightly better under ML profiles, and while this performance drops for Hard instances, these models almost never return invalid or partial matchings.
Prompting Techniques. Prompt-engineering techniques have been empirically demonstrated to enhance the performance of LLMs on mathematical reasoning and formal logic inference tasks [wei2022chain, Yao2023ToT, Besta2024GoT]. We evaluated a range of prompt-engineering techniques—including few-shot prompting and Chain-of-Thought (CoT) prompting, which supply exemplar “thought processes” and intermediate reasoning steps—in an attempt to bolster LLM performance. However, none of these strategies qualitatively improved on medium- or hard-difficulty instances. See Appendix˜D for further details.
@@ 3.1 Measuring Instability
A natural question is how far LLM‐produced responses deviate from stable outcomes. To quantify this, we use two complementary metrics: instability rate and optimality rates (see Section˜2.2). The instability rate directly reflects the distance from any stable solution, whereas the optimality rate implicitly evaluates the model’s success in executing the underlying matching procedure. Figure˜4 and Figure˜4 illustrate comparisons of LLMs with a baseline of randomly generated outcomes. [^5] Note that the plots only illustrate unstable but valid one-to-one outputs.
@@
Interestingly, the performance distinction between basic and reasoning models becomes less clear. While the intermediate reasoning models return a lower instability rate in Easy and Medium problems, their performance significantly drops in larger-scale problems (Hard). In fact, the performance of DeepSeek-70B becomes worse than even basic non-reasoning models. We attribute this behavior to the model’s diminished capacity for handling larger input lengths, a hypothesis supported by their lower proportion of valid outcomes (as seen in Figure˜2).
@@ 4 Resolving Instability
Generating stable solutions requires both exact reasoning over agents’ preference lists and the execution of a stability-guaranteeing procedure (e.g., the DA algorithm). As demonstrated in Section˜3, all evaluated models—irrespective of their reasoning capabilities—exhibit severe performance degradation as the problem size grows. This leads to the natural question of whether these models can resolve instability in a given matching—a task that entails detecting blocking pairs through preference reasoning and applying an appropriate sequence of adjustments to restore stability.
We provide LLMs with unstable (but valid) matchings along with preference profiles, and instruct them to convert these initialized solutions to stable matchings. To assess how the instability rate may influence solution quality, we distinguish two classes of initial matchings: (i) One-BP, matchings containing exactly one blocking pair (i.e., “almost stable”) such that their stability may be resolved through simpler proposal-rejection iterations, and (ii) Random, matchings sampled uniformly at random from the set of all valid one‐to‐one pairings, which typically contain a large number of blocking pairs and thus exhibit high degrees of instability. See Appendix˜F for detailed steps and pseudo-code for generating one-BP and random initialization. Note that starting from an arbitrary matching, sequentially resolving blocking pairs may result in a cycle—as shown by knuth1976marriages. However, a random sequence converges to stability with probability one [roth1990random, abeledo1995paths].
@@
Table˜1 displays the fraction of responses in which LLMs return stable matchings when asked to resolve the above types of unstable matching. Surprisingly, our experiments illustrate that in the task of resolving instability, the performance of all evaluated models does not exceed—and even degrades—their performance in generating stable solutions. This behavior persists regardless of initial matchings (One-BP or Random) and LLMs’ reasoning capability. In fact, on Hard instances, the output returned by advanced reasoning models on One-BP matchings (i.e., containing a single blocking pair) contains substantially more than one blocking pair. In other words, even for the most basic instances, LLMs often introduce additional instabilities beyond the original single violation. We elaborate on this in Appendix˜F, demonstrating how all models, including those with advanced reasoning, often return solutions with a higher instability rate, highlighting their inability to systematically eliminate blocking pairs in accordance with preference lists.
@@ 5 Detecting Instability
The findings in previous sections raise the question of whether LLMs can reliably detect instability in a given matching—a simpler task that involves iterating over each unmatched pair to determine whether both agents prefer one another over their assigned partners. This procedure entails only a straightforward comparison of preferences and requires $\mathcal{O}(n^{2})$ steps.
@@
Reasoning Models: Preferences and Blocking Pairs. Figure˜5 reveals an interesting observation about reasoning models: their performance is influenced by the number of blocking pairs in the matching being evaluated—similar to the observations in Section˜4. They achieve a high accuracy on identifying random matchings (which have a larger number of blocking pairs) as unstable and a significantly lower accuracy with matchings that have at most one blocking pair.
Basic Models and Hallucination. Interestingly, the non-reasoning models achieve a high accuracy ($80\%$) with both types of unstable matchings, and extremely low ($20\%$) accuracy with stable matchings. Note that the performance across all models is qualitatively similar in ML and IC profiles, even though each ML profile admits a unique stable solution (thus, identical reports for Optimal and Pessimal). See Appendix˜G for further analysis. A manual analysis of non‐reasoning models uncovers frequent hallucinations about blocking pairs, resulting in a systematic bias toward classifying matchings as unstable. This can be largely attributed to misinterpretations of the input preferences.
@@ 6 Reasoning about Ranked Preferences
Many advanced reasoning paradigms, ranging from causal inference [Haoang2024Causal] and counterfactual analysis to game-theoretic decision making, depend fundamentally on the ability to compare and evaluate alternative choices. As demonstrated thus far, even advanced reasoning models often fail to execute the step-by-step procedures of combinatorial algorithms when those procedures operate over ranked preference lists. This shortcoming motivates the question of whether current LLMs can truly reason about preferences, as opposed to merely applying preferences in generating responses heuristically.
@@
Hierarchical, level‐wise reasoning evaluations have been proposed recently in domains such as causal inference of LLMs [Haoang2024Causal]. For example, an L1 question is “Who is agent W5’s, 4th-most preferred agent?”, and an L2 question “Would agent W5, prefer M8 over M7?”
@@
Basic models have low accuracy in all levels of difficulty even on small instances, which is probably the reason behind their inability to compute or detect stable solutions (as discussed in Section˜3 and Section˜5). In basic models (e.g., Llama-3.3-70B) and (non-advanced) reasoning models (e.g., DeepSeek-70B), the size of the problem has a greater impact on accuracy as compared to the level of the question, indicating their difficulty of handling larger inputs. Although advanced-reasoning models have significantly higher accuracy rates compared to other models, they still make minor errors, especially in larger profiles (Hard). Generating stable solutions for these instances often requires a larger number of reasoning steps over many preference lists, causing minor errors to compound (small errors multiply!).
@@ 7 Performance Improvement through LoRA Fine-Tuning
Supervised fine-tuning has proven to be effective in enhancing the reasoning capabilities of LLMs in a variety of tasks, including mathematical problem solving [Chung2024Flan, Lewkowycz2022Quant], logical reasoning [Morishita2023Logical], and code generation [Luo2024Wizard]. Markeeva2024CLRS demonstrate how fine-tuning a small LLM (2B parameters) can significantly improve LLMs’ ability to execute textbook algorithms (e.g., sorting an array, finding the shortest path between two nodes on a graph, etc.). Hence, we evaluate whether fine-tuning can be used to enhance LLMs’ ability to compute stable matchings—a task requiring reasoning over ranked preferences in addition to algorithmic understanding. To this end, we perform LoRA fine-tuning [Hu2022LoRA] on three reasoning models, including Qwen-QwQ-32B and two smaller models, DeepSeek-8B and DeepSeek-14B, for Generating Stable Solutions task. Additionally, we also evaluate whether fine-tuning can mitigate errors made by LLMs in the Preference Reasoning task (Section˜6).
Training. Let $\mathcal{D}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N}$ be the fine-tuning dataset for a given task. Each pair consists of an input instance $\mathbf{x}^{(i)}$ and the desired model completion $\mathbf{y}^{(i)}$. The input instance $\mathbf{x}^{(i)}$ is made up of four components: (i) a generic system-prompt $s$, (ii) a high-level instruction $u$, (iii) the preference profile $p^{(i)}$, and (iv) the task-prompt $t^{(i)}$. The desired completion $\mathbf{y}^{(i)}$ consists of two components, (i) a chain-of-thought reasoning trace $r^{(i)}$, and (ii) the answer $a^{(i)}$ in the desired format. Each model is fine-tuned with standard next-token cross-entropy on the concatenated sequence $\mathbf{z}^{(i)}=\mathbf{x}^{(i)}||\mathbf{y}^{(i)}$. We separately fine-tune each model for the Generating Stable solutions task ($N=10,000$) and the Preference Reasoning task ($N=9,000$). See Appendix˜H for details of the fine-tuning setup and results from experiments with different hyperparameter values.
Improvement. Fine-tuning LLMs with data containing synthetically generated reasoning traces substantially improves their performance on both tasks, as evidenced in Table˜2.
@@
In fact, this approach enhances the performance of Qwen-QwQ-32B, a (non-advanced) reasoning model, to a success rate of $100\%$ in computing stable matchings with both Easy and Medium instances, significantly outperforming advanced-reasoning models. Fine-tuning also clearly improves smaller models, i.e. DeepSeek-8B and DeepSeek-14B, both in terms of achieving stability and Instability Rate. [^6] Interestingly, this improvement is clearer for ML instances, where both models achieve a $100\%$ success rate at the Easy level and $80\%$ success rate at the Medium level. Similar results are obtained for the Preference Reasoning tasks, with the error-rate reducing to 0 at the Easy and Medium levels.  [^7] The only exception being DeepSeek-8B at the Medium level.
@@
Category Model Stable Solutions ($\%$) Instability Rate ($\downarrow$) Optimality Rate ($\uparrow$) Stable Solutions ($\%$) Instability Rate ($\downarrow$) Optimality Rate ($\uparrow$) Accuracy ($\%$) Accuracy ($\%$)
@@ 8 Concluding Remarks
We summarize the performance of LLMs across all four tasks in Table˜3, reflecting the clear hierarchy between advanced-reasoning models, (non-advanced) reasoning models, and basic models. The limitations in reliably reasoning about ranked preferences raise concerns about the viability of LLMs as agents acting on behalf of users in market-oriented or preference-sensitive decision-making settings, limit their capacity to negotiate complex user preferences, and hinder efforts in developing pluralistic techniques (e.g., constitutional AI [bai2022constitutional] and social choice-theoretic [Conitzer+24]) for value alignment that are inherently based on aggregating rankings.
Open-Source vs. Closed-Source models. Among the models that we evaluate, Gemini-2.5-Pro (a closed-source model) emerges as the most capable across all tasks. While DeepSeek-R1 (open-source) broadly outperforms OpenAI’s o3-mini (closed-source), it performs much worse with IC instances than with ML instances. While both basic models struggle on all tasks, Gemini-2.0-Flash (closed-source) marginally outperforms Llama-3.3-70B (open-source) on various metrics. Given the promising improvement yielded by fine-tuning an open-source reasoning model, i.e. Qwen-QwQ-32B, it is worth exploring strategies that also enable it to handle large inputs.
Beyond Linear Preferences. Our current evaluation paradigm considers complete and strict linear preferences. In real-world scenarios, however, preferences involve complexities such as incompleteness, indifference between alternatives, and capacity constraints [Delorme2019Incomplete, Kawanashie2013Hospital, Manlove2O22Student]. As a result, algorithms to compute stable solutions in such settings are far more complex and solutions are often intractable [Manlove2014Book]. While, in Appendix˜E, we provide some preliminary insights on preferences with ties, understanding such cases requires deeper theoretical and empirical investigation. A meaningful next step would be to examine how AI models respond to these more intricate preference structures.